# Online Feature Updates Improve Online (Generalized) Label Shift Adaptation

### v1

@@ 3.3 Online Feature Updates Improve Online Generalized Label Shift Adaptation
When we have knowledge of the feature map $h$ under generalized label shift, the problem simplifies to the classical online label shift scenario. However, the more challenging situation arises when $h$ remains unknown. Due to the violation of the label shift assumption, standard OLS methods can perform arbitrarily bad. Fortunately, existing research in test-time training (TTT) (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022) demonstrates that feature updates driven by SSL align the source and target domains in feature space. When the source and target domains achieve perfect alignment, such feature extractor effectively serves as the feature map $h$ as assumed in generalized label shift. Therefore, the sequence of feature extractors in $f_{1},\cdots,f_{T}$ generated by Algorithm 1 progressively approximates the underlying $h$. This suggests that, compared to the original OLS, OLS with online feature updates (Algorithm 1) experiences a milder violation of the label shift assumption within the feature space and is actually expected to have better performance in the setting of online generalized label shift. Indeed, as demonstrated in Section 4, we can observe significant improvements in OLS-OFU compared to standard OLS for online generalized label shift.
@@ 4 Experiment
In this section, we empirically evaluate how OLS-OFU improves the original OLS methods on both online label shift and online generalized label shift [^2] Code released at https://anonymous.4open.science/r/online_label_shift_with_online_feature_updates-3A1F/.. The experiment is performed on CIFAR-10 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), EuroSAT (Helber et al., 2019), and CIFAR-10C (Hendrycks & Dietterich, 2019). We systematically vary the distribution shift processes and self-supervised learning (SSL) techniques to assess the effectiveness of the proposed method.
@@ 4.1 Experiment Set-up
Dataset and Label Shift Settings. For online label shift, we evaluate the efficacy of our algorithm on CIFAR-10 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), and EuroSAT (Helber et al., 2019), which have 10 categories of images. For each dataset, we split the original train set into the offline train (i.e., $D_{0}$) and validation sets (i.e., $D^{\prime}_{0}$) following a ratio of $4:1$. At the online test unlabeled batches are sampled from the test set. For online generalized label shift, the offline train and validation sets are the same CIFAR-10 images. The test unlabeled batches are drawn from CIFAR-10C, a benchmark with the same objects as CIFAR-10 but with various types of corruption. We experiment with three types of corruptions with CIFAR-10C: Gaussian noise, Fog and Pixelate.
We follow Bai et al. (2022) and Baby et al. (2023) to simulate the online label distribution shift with two shift patterns: Sinusoidal shift and Bernoulli shift. Specifically, given two label distribution vectors $q$ and $q^{\prime}$, we simulate the label marginal distributions at time $t$ as a weighted combination of them: $q_{t}:=\alpha_{t}q+(1-\alpha_{t})q^{\prime}$. In Sinusoidal shift, $\alpha_{t}=\sin\frac{i\pi}{L}$ (periodic length $L=\sqrt{T}$, $i=t\textnormal{ mod }L$) while in Bernoulli shift, $\alpha_{t}$ is a random bit (either 0 or 1), where the bit switches $\alpha_{t}=1-\alpha_{t-1}$ with probability $p=1-\frac{1}{\sqrt{T}}$. In our experiments, we set the initial distribution vector $q$ and $q^{\prime}$ with $\frac{1}{K}(1,\cdots,1)$ and $(1,0,\cdots,0)$. To sample the batch test data at time $t$, we first sample a batch of labels (not revealed to the learner) according to $q_{t}$. Then given each label we can sample an image from the test set, and collect this batch of images without labels as $S_{t}$. We experiment with $T=1000$ and batch size $B=10$ at each time step, following Baby et al. (2023).
Evaluation Metric. We report the average error during test, i.e., $\frac{1}{TB}\sum_{t=1}^{T}\sum_{x_{t}\in S_{t}}\mathbbm{1}\left(f_{t}(x_{t})\neq y_{t}\right)$, where $(x_{t},y_{t})\sim\mathcal{P}_{t}^{\rm test}$, to approximate $\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t};\mathcal{P}^{\rm test}_{t})$ for the evaluation efficiency. This is valid for large $T$ due to its exponential concentration rate by the Azuma–Hoeffding inequality.
Baselines and OLS Methods. We perform an extensive evaluation of 6 OLS methods in the literature: FTFWH, FTH, ROGD, UOGD, ATLAS, and FLHFTL. We further report the performance of our method OLS-OFU (Algorithm 1) applied on top of each OLS. OLS-OFU is instantiated with 3 common SSL methods introduced in Section 2: rotation degree prediction, entropy minimization, and MoCo. We run the batch accumulation variant of Algorithm 1 for MoCo and set BA=$50$; please see the ablation for the selection of BA in Appendix E.8.1. Additionally, we report two baseline scores, Base and OFU. The Base uses the fixed pre-trained model $f_{0}$ to predict the labels at all time steps. The second is online feature updates (OFU) only, where at time step $t$ we only update the features (Step 2 in Algorithm 1) without utilizing OLS algorithms.
@@ 4.2 Results
Main Results: comparison between OLS-OFU and baselines under online (generalized) label shift. Figure 2(a) shows the performance comparison between OLS-OFU, implemented with three SSL methods, and their corresponding OLS counterparts on CIFAR-10 under the scenario of classical online label shift. Figure 2(b) shows the results on three more datasets with SSL technique in OLS-OFU being rotation degree prediction. The online shift pattern in the setting of Figure 2(a-b) is sinusoidal shift. From Figure 2(a-b), we can observe that OLS-OFU outperforms all other baselines. Notably, OLS-OFU, when integrated with three SSL methods, consistently outperforms their counterpart OLS method across all of the OLS methods. This emphasizes the importance of improving representation learning within label shift adaptation and underscores the robustness of our framework OLS-OFU. Results for Bernoulli shift in Appendix E.3 show similar conclusion.
Furthermore, we report the performance of OLS-OFU in the context of online generalized label shift. In this scenario, the test images exhibit three types of domain shifts in CIFAR-10C — Gaussian noise, Fog, and Pixelation—with mild severity. We present the result under Sinusoidal shift with the SSL method used in OLS-OFU as rotation degree prediction. Similar to the scenario of online label shift, the superior performance of OLS-OFU compared with Base, OFU, and OLS demonstrates its superiority (Figure 2(c)). Particularly, OLS-OFU consistently outperforms the original OLS methods by an even larger gap than the one occurring at the classic label shift setting. We hypothesize that the inclusion of SSL methods within OLS-OFU aids the learning of the unknown mapping $h$ inherent in the generalized label shift assumption. As evaluated in Sun et al. (2020), the label shift assumption in the feature space of $\theta^{\rm feat}_{t}$ is violated milder than the one of $\theta^{\rm feat}_{0}$ — the more severe violation of $\theta^{\rm feat}_{0}$ explains why some OLS methods perform even worse than OFU. Results for other SSL methods and online shift patterns show a similar pattern in Appendix E.4.
Moreover, as an ablation study of OLS-OFU, the comparison between OLS-OFU and either OLS or OFU individually underscores the specific benefits of both the OFU and OLS modules in addressing the challenges posed by the online (generalized) label shift problem.
@@
Empirical validation of Equation 8. In Section 3.2, we argued that when the inequality in Equation 8 holds, the loss of FLHFTL-OFU exhibits a tighter upper bound compared to FLHFTL. Figure 3 presents the RHS (corresponds to OLS) and LHS (corresponds to OLS-OFU with SSL loss as rotation degree prediction) of Equation 8. We perform the study over eight different settings and varies the domain shift and online shift patterns. It is evident that OLS-OFU yields improvements on the baseline of the regret as shown in Equation 8. Appendix E.6 validates this inequality for other SSL techniques. We further examine how the improvement from Equation 8 (RHS - LHS) is correlated the improvement from OLS to OLS-OFU. For each OLS method, we compute the Pearson coefficient (within the range of [-1, 1]) between these two terms across all experimental setups (varying dataset, online shift pattern, and SSL methods). The calculated coefficients are 0.70, 0.76, 0.54, 0.63, 0.58, and 0.75 when OLS is FTFWH, FTH, ROGD, ATLAS, UOGD, or FLHFTL, respectively. These values indicate strong positive correlations, underscoring how improvements in regret align with empirical performance improvements.
Results on CIFAR-10C with higher severity. We evaluate all methods on CIFAR-10C with higher severity of domain shifts. Figure 4 shows the results when the test images are corrupted by Gaussian noise [^3] Please refer to Appendix E.5 for results involving higher levels of severity, other corruption types, and SSL techniques.. Comparing with Figure 2(c), it is clear that as the severity increases, OLS-OFU enlarges the performance gap compared to OLS. However, it might be worth pointing out that neither OLS nor OLS-OFU are better than OFU.
Figure 4: Results on CIFAR-10C for a high level of severity.
[…]
5 Conclusion
@@
We focus on online (generalized) label shift adaptation and present a novel framework Online Label Shift adaptation with Online Feature Updates (OLS-OFU), which harnesses the power of self-supervised learning to enhance feature representations dynamically during testing, leading to improved predictive models and better test time performance.
@@
Discussion and future work. One promising direction is to extend the idea of this paper to online covariate shift, for example, the algorithm in Zhang et al. (2023) freezes the feature extractor and only updates the linear layer. Another possible direction is to consider a more realistic domain shift scenario within the generalized label shift setting — domain shift types may vary over time or be even more challenging, such as shifting from cartoon images to realistic images.
Acknowledgements
@@
RW and KQW are supported by grants from LinkedIn, DARPA Geometries of Learning, and the National Science Foundation NSF (1934714). RW is also supported by grants from the National Science Foundation NSF (2241100), Army Research Office ARO MURI (W911NF2110317), and Office of Naval Research ONR (N00014-20-1-2334).
@@ Appendix A Further Related Work
Offline distribution shift and domain shift. Offline label shift and covariate shift have been studied for many years. Some early work (Saerens et al., 2002; Lin et al., 2002) assumes the knowledge of how the distribution is shifted. Later work (Shimodaira, 2000; Zadrozny, 2004; Huang et al., 2006; Gretton et al., 2009; Lipton et al., 2018; Alexandari et al., 2020; Azizzadenesheli et al., 2019; Garg et al., 2020) relaxes this assumption and estimates this knowledge from unlabeled test data. A recent work by (Garg et al., 2023) considers a relaxed version of offline label shift problem where the class-conditionals can change between train and test domains in a restrictive way. Extending their results to the case of relaxed online version of generalized label shift is an interesting future direction.
Online distribution shift with provable guarantees. There has been several work modeling online distribution shift as the classic online learning problem (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023; Zhang et al., 2023), which leverage the classical online learning algorithms (Shalev-Shwartz, 2012; Besbes et al., 2015; Baby & Wang, 2022) to bound the static or dynamic regret. However, none of them updates the feature extractor in a deep learning model but only the last linear layer or the post-hoc linear reweighting vectors. Our proposed method OLS-OFU utilizes the deep learning SSL to improve the feature extractor, which brings better performance.
Domain shift adaptation within online streaming data. When we consider the most authentic online learning setup where the learner only receives the unlabeled samples, the most representative idea is test-time training (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022), which utilizes a (deep learning) self-supervised loss to online update the model. However, it focuses on how to adapt to a fixed domain shifted distribution from online streaming data and is not designed for how to adapt to continuous distribution changes during the test stage, while our algorithm concentrates the later problem. Besides test-time training, Hoffman et al. (2014) and (Mullapudi et al., 2019) study the online domain shift for specific visual applications.
Algorithm 3 Revised ROGD for online feature updates ROGD-R. See the original version in Equation 7 and Equation 8 in (Wu et al., 2021).
@@
Algorithm 4 Revised FTH for online feature updates (FTH-R). See the original version in Equation 9 in (Wu et al., 2021).

### v3

@@ 3.4 Online Feature Updates Improve Online Generalized Label Shift Adaptation
The generalized label shift is harder because the feature map $h$ is unknown and naively applying OLS methods might raise the challenge due to the violation of the label shift assumption. Fortunately, existing research in test-time training (TTT) [46, 48, 36, 39] demonstrates that feature updates driven by SSL align the source and target domains in feature space. When the source and target domains achieve perfect alignment, such feature extractor effectively serves as the feature map $h$ as assumed in generalized label shift. Therefore, the sequence of feature extractors in $f_{1},\cdots,f_{T}$ generated by Algorithm 1 progressively approximates the underlying $h$. This suggests that, compared to the original OLS, OLS-OFU experiences a milder violation of the label shift assumption within the feature space and is expected to have better performance in online generalized label shift settings.
@@ 4 Experiment
FTFWH FTH ROGD ATLAS UOGD FLH-FTL
OLS 11.9% / 0.441 12.04% / 0.705 13.65% / 0.937 12.18% / 43.2 11.54% / 297 12.02% / 1.15
OLS-OFU ($\tau=1$) 11.3% / 283 11.2% / 286 13.9% / 385 11.6% / 47 11.4% / 319 11.2% / 285
OLS-OFU ($\tau=10$) 9.93% / 28.8 10.2% / 29.6 13.1% / 41.9 11.6% / 45.1 10.5% / 306 9.93% / 29.9
OLS-OFU ($\tau=50$) 8.15% / 6.47 8.89% / 6.91 12% / 9.49 11.3% / 44.8 9.27% / 303 8.27% / 7.08
OLS-OFU ($\tau=100$) 7.33% / 3.57 8% / 3.85 10.8% / 5.29 10.1% / 44.2 8.35% / 303 7.45% / 4.15
OLS-OFU ($\tau=500$) 11.6% / 1.27 11.7% / 1.53 13.5% / 2.09 12.1% / 43.6 11.3% / 298 11.7% / 1.94
Table 1: Average error/time (minutes) of 6 original OLS methods versus OLS-OFU with various frequency $\tau$ in batch accumulation. The SSL in OLS-OFU is the rotation degree prediction.
In this section, we initiate OLS-OFU with three popular SSL techniques and empirically evaluate how OLS-OFU improves the original OLS methods on both online label shift and online generalized label shift on various datasets and shift patterns [^1] Code is released at https://github.com/dattasiddhartha/online-feature-updates-olsofu.
@@ 4.1 Experiment Set-up
Dataset and Label Shift Settings. For online label shift, we evaluate the efficacy of our algorithm on CIFAR-10 [29], STL10 [12], CINIC [13], and EuroSAT [25]. For each dataset, we split the original train set into the offline train (i.e., $D_{0}$) and validation sets (i.e., $D^{\prime}_{0}$) following a ratio of $4:1$. At the online test stage, unlabeled batches are sampled from the test set. For online generalized label shift, the offline train and validation sets are the CIFAR-10 images. The test unlabeled batches are drawn from CIFAR-10C [26], a benchmark with the same objects as CIFAR-10 but with various types of corruption. We experiment with three types of corruptions with CIFAR-10C: Gaussian noise, Fog, and Pixelate. We follow Bai et al. [7] and Baby et al. [6] to simulate the online label distribution shift with two shift patterns: Sinusoidal shift and Bernoulli shift. We experiment with $T=1000$ and batch size $B=10$ at each time step, following Baby et al. [6]. See more details of dataset set-up and online shift patterns in Appendix E.1.
Evaluation Metric. We report the average error during test, i.e., $\frac{1}{TB}\sum_{t=1}^{T}\sum_{x_{t}\in S_{t}}\mathbbm{1}\left(f_{t}(x_{t})\neq y_{t}\right)$, where $(x_{t},y_{t})\sim\mathcal{P}_{t}^{\rm test}$, to approximate $\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t};\mathcal{P}^{\rm test}_{t})$ for the evaluation efficiency.
Self-supervised learning methods in OLS-OFU. In the experiment, we narrow our focus on three particular SSL techniques in the evaluation for classification tasks: rotation degree prediction [16, 46], entropy minimization [18, 48] and MoCo [23, 10, 11]. It is important to note that this concept extends beyond these three SSL techniques, and the incorporation of more advanced SSL techniques to further elevate the performance. Appendix E.1 gives more details of these SSL techniques.
Set-ups of OLS methods, OLS-OFU and baselines. We perform an extensive evaluation of 6 OLS methods in the literature: FTFWH, FTH, ROGD, UOGD, ATLAS, and FLHFTL by following the implementation in Baby et al. [6]. We report the performance of our method OLS-OFU (Algorithm 1) applied on top of each OLS and 3 SSL methods introduced above. The frequency parameter $\tau$ is fixed as $100$ for most experiments unless we particularly mention it. Additionally, by following the setup in [49, 6], we report one baseline score Base, which uses the fixed pre-trained model $f_{0}$ to predict the labels at all time steps.
@@ 4.2 Results
Main Results: comparison between OLS-OFU and OLS under online (generalized) label shift. Figure 2(a) shows the performance comparison between OLS-OFU, implemented with three SSL methods, and their corresponding OLS counterparts on CIFAR-10 under the scenario of classical online label shift. Figure 2(b) shows the results on three more datasets with SSL technique in OLS-OFU being rotation degree prediction. Figure 2(c) shows the results on CIFAR-10C datasets for evaluating methods on online generalized label shift. The online shift pattern in Figure 2 is the sinusoidal shift and similar results on the Bernoulli shift are in Appendix E.2. We have two main observations from the results. First, we find our OLS-OFU method achieves substantial improvements over existing OLS methods, which is as significant as to the gains existing OLS methods have over the baseline (i.e., without distribution shift adaptations). This demonstrates that integrating online feature updates is as effective in solving online distribution shifts as the fundamental online label shift method itself. Second, the improvement is consistent across all six original OLS methods and three SSL techniques on all datasets, which further demonstrates our OLS-OFU is general enough to incorporate future OLS methods and more advanced SSL techniques as well.
Validating Principle 1: ablation study on the order between OLS and the feature update step in OLS-OFU. We compare OLS-OFU with its other variant named OLS-OFU-difforder where we update SSL first and run OLS later (which violates Principle 1). We compare these two algorithms across all previous 6 OLS methods and two choices of batch accumulation $\tau=1$ and $\tau=100$. The dataset is CIFAR-10, the SSL is rotation degree prediction and the shift pattern is the sinusoidal shift. We report the average error every $\tau$ time step when the feature extractor is updated, so the order matters. From Table 2, we can observe that benefiting from Principle 1, the error OLS-OFU is consistently lower than OLS-OFU-difforder. This means that Principle 1 indeed is not only necessary for the correctness of theoretical guarantee but also plays an important role in practice.
FTFWH FTH ROGD ATLAS UOGD FLHFTL
OLS-OFU($\tau=1$) 11.3% 11.2% 13.9% 11.6% 11.4% 11.2%
OLS-OFU-difforder($\tau=1$) 12.33% 12.12% 14.35% 12.10% 11.91% 12.08%
OLS-OFU($\tau=100$) 7.23% 7.91% 10.81% 10.12% 8.33% 7.53%
OLS-OFU-difforder($\tau=100$) 8.02% 8.63% 11.19% 10.55% 8.80% 8.29%
Table 2: Ablation study on the order between OLS and the feature update step in OLS-OFU.
Validating Principle 3: ablation study on frequency $\tau$ in batch accumulation in OLS-OFU. In section 3.1, we introduce the batch accumulation to flavor the effectiveness of the SSL and reduce the additional time cost of OLS-OFU compared with the original OLS methods. In Table 1, we evaluate OLS-OFU on CIFAR-10 with various $\tau$ while initiating OLS-OFU with rotation degree prediction and compare the average error and time with OLS. We have these observations across all OLS (columns): 1. OLS-OFU with all $\tau$ outperforms OLS; 2. The average error of OLS-OFU decreases from $\tau=1$ (w/o batch accumulation) to $\tau=100$ and starts to increase after and this is because the effectiveness of SSL benefits from $\tau>1$ while larger $\tau$ means less updates and hence hurts the long term performance; 3. the additional time cost of OLS-OFU with larger $\tau$ is smaller and $\tau=100$ gives a great balance of performance and time cost for all OLS (and SSL; the similar tables for other two SSL are presented in the Appendix E.3). From the results, we recommend $\tau=100$ as a good starting point for choosing this parameter.
Ablation study on the amount of stored training data in step (3). We further study how the amount of training data stored influences the effectiveness of our OLS-OFU when the OLS is ROGD, FTH, or FLHFTL, which is dependent on step (3). We experiment with the CIFAR-10 dataset and the shift pattern of sinusoidal shift; the SSL is chosen as rotation degree prediction. The percentage of stored training data (out of the whole training set of size 10,000) is varied in $\{100\%,80\%,60\%,40\%,20\%,10\%,5\%,0\%\}$; $0\%$ means that we still update the feature extractor but reuse the pretrained linear classifier. The results are reported in Table 3.
Figure 3: Empirical validation of Equation 7. Clean denotes the experiment on CIFAR-10, and others denote the corruption type. They are paired with two online shift patterns: Sinusoidal and Bernoulli.
We can observe that with less stored training data for retraining the last linear layer, the error of OLS-OFU would increase gradually. However, an important finding is that even with $0\%$ stored training data, the error of OLS-OFU is still lower than the OLS without feature extractor updates. This can actually explained by the original test-time training papers [46, 48, 36, 39], where they only update the feature extractor without refining the last linear layer and still have substantial benefit. The results suggest that a large amount of stored training data is not necessary for the effectiveness of OLS-OFU, while more stored data can bring more benefits.
% Stored Training Data 100% 80% 60% 40% 20% 10% 5% 0% OLS only
FTH-OFU 8% 8.18% 8.68% 9.49% 9.54% 9.67% 9.81% 10.24% 12.04%
ROGD-OFU 10.8% 10.93% 11.84% 12.50% 12.63% 12.82% 12.94% 13.50% 13.65%
FLHFTL-OFU 7.45% 7.62% 8.04% 8.91% 9.51% 10.11% 10.34% 10.43% 12.02%
FTFWH-OFU 7.33% 7.48% 7.92% 8.40% 8.91% 9.86% 10.03% 10.41% 11.9%
Table 3: The effectiveness of OLS-OFU versus the percentage of stored training data in step (3).
@@
Empirical validation of Equation 7. In Section 3.3, we argued that when the inequality in Equation 7 holds, the loss of FLHFTL-OFU exhibits a tighter upper bound compared to FLHFTL. Figure 3 presents the RHS (corresponds to OLS) and LHS (corresponds to OLS-OFU with SSL loss as rotation degree prediction) of Equation 7. We perform the study over eight different settings and vary the domain shift and online shift patterns. It is evident that OLS-OFU yields improvements on the baseline of the regret as shown in Equation 7. Appendix E.4 validates this inequality for other SSL.
5 Conclusion and Future Work
@@
We focus on online (generalized) label shift adaptation and present a novel framework OLS-OFU, which harnesses the power of self-supervised learning to enhance feature representations dynamically during testing, leading to improved predictive models and better test time performance.
@@
Discussion and future work. One promising direction is to extend the idea of this paper to online covariate shift, for example, the algorithm in Zhang et al. [53] freezes the feature extractor and only updates the linear layer. Another possible direction is to consider a more realistic domain shift scenario within the generalized label shift setting — domain shift types may vary over time or be even more challenging, such as shifting from cartoon images to realistic images.
Acknowledgments and Disclosure of Funding
@@
RW and KQW are supported by grants from LinkedIn, DARPA Geometries of Learning, and the National Science Foundation NSF (1934714). RW is also supported by grants from the National Science Foundation NSF (CIF-2402817, CNS-1804829), SaTC-2241100, CCF-2217058, ARO-MURI (W911NF2110317), and ONR under N00014-24-1-2304. YW and DB are partially supported by NSF Award #2134214.
@@ Appendix A Further Related Work
Offline distribution shift and domain shift. Offline label shift and covariate shift have been studied for many years. Some early work [42, 33] assumes the knowledge of how the distribution is shifted. Later work [45, 51, 28, 19, 35, 1, 4, 14] relaxes this assumption and estimates this knowledge from unlabeled test data. A recent work by [15] considers a relaxed version of offline label shift problem where the class-conditionals can change between train and test domains in a restrictive way. Extending their results to the case of relaxed online version of generalized label shift is an interesting future direction.
Online distribution shift with provable guarantees. There has been several work modeling online distribution shift as the classic online learning problem [49, 7, 6, 53], which leverage the classical online learning algorithms [44, 8, 5] to bound the static or dynamic regret. Qian et al. [40] focuses on a special setting in online label shift, where there can be new classes occurring in the test stage. They handle the new class by unsupervised estimating the portion of unseen data. However, none of them updates the feature extractor in a deep learning model but only the last linear layer or the post-hoc linear reweighting vectors. Our proposed method OLS-OFU utilizes the deep learning SSL to improve the feature extractor, which brings better performance.
Domain shift adaptation within online streaming data. When we consider the most authentic online learning setup where the learner only receives the unlabeled samples, the most representative idea is test-time training [46, 48, 36, 39], which utilizes a (deep learning) self-supervised loss to online update the model. However, it focuses on how to adapt to a fixed domain shifted distribution from online streaming data and is not designed for how to adapt to continuous distribution changes during the test stage, while our algorithm concentrates the later problem. Besides test-time training, Hoffman et al. [27] and [38] study the online domain shift for specific visual applications.
Algorithm 2 Revised ROGD for online feature updates ROGD-R. See the original version in Equation 7 and Equation 8 in [49].
@@
Algorithm 3 Revised FTH for online feature updates (FTH-R). See the original version in Equation 9 in [49].